OpenStudy (anonymous):

would u know where these 2 linear equations intersects? if it does?....2x-y=-2 and y=2x+4

OpenStudy (anonymous):

2x - y = -2 <=> y = 2x + 2 vs. y = 2x + 4 You have two parallel lines intersecting the y-axis at two different points, so no they will not intersect.

OpenStudy (anonymous):

can you please how did u arrive to that solution? do i need to know what is x and y?

OpenStudy (anonymous):

It would be easier if you just plotted the two lines by hand. Start by plugging in x = -5 into both equations and checking the resulting y value. Then do x = -4, -3, ..., 0, 1, 2, ...5. You'll see that they are parallel. Also, both lines have the same slope ( the ratio of y to x ) so they will be parallel.

OpenStudy (anonymous):

You can get slope by looking at the a in the equation y = ax + b. The slope of those two lines is a = 2.

OpenStudy (anonymous):

so meaning to say, i dont need to compute and find the equivalent of x and y?

OpenStudy (anonymous):

i dont really understand...would you mind..i still have 4 more? can u help me?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

I'll take a look at one more, sure

OpenStudy (anonymous):

x-5y=15 and x+ 5y=-5

OpenStudy (anonymous):

would u mind doing it step by step for me to follow you? :) pleasee

OpenStudy (anonymous):

Ok, solving for the first equation you see that x - 5y = 15 <=> 5y = x - 15 Plug this into the second equation x + 5y = -5 You get x + ( x - 15 ) = -5 Simplify and you get 2x = 10, x = 5 Plug x = 5 into either of the equations to get y = - 2

OpenStudy (anonymous):

Therefore, the two lines intersect at x = 5, y = 2 ( 5, - 2 )

OpenStudy (anonymous):

what the pairs?

OpenStudy (anonymous):

are you still therE?

OpenStudy (anonymous):

Yes, what do you mean?

OpenStudy (anonymous):

pairs like for me to get where it intersects...

OpenStudy (anonymous):

There is only one point at which it intersects.

OpenStudy (anonymous):

Two straight lines can only intersect at one point on a flat 2D surface.

OpenStudy (anonymous):

ok..i will just review what you just typed later...can you still answer 2 more?

OpenStudy (anonymous):

2x-3y=6 and 6x=18+9y

OpenStudy (anonymous):

just so amazed by your answers...

OpenStudy (anonymous):

thank you for helping me with the first 2

OpenStudy (anonymous):

still around?

OpenStudy (anonymous):

hello...?